An ac generator has a frequency of 5.5 kHz and a voltage of 30 V. When an inductor is connected between the terminals of this generator, the current in the inductor is 30 mA. What is the inductance of the inductor? L= ?
Answer #1

Given that Ac generator frequency = 5.5Khz

Voltage = 30 volts.

Now current through the inductor = 30mA

For the case given, current through the inductor will be = V/(L) where L is the inductive reactance.

As = 2*pi*f

Thus I = V/(2*pi*f*L)

So L = V/(2*pi*f*I)

L = 30/(2*pi*5.5*10^3*30*10^(-3))

L = 30/ 1036.72 = 0.0289 Henry

L = 0.0289 Henry = 28.9 mH. Ans